OpenStudy (anonymous):

so someone help me figure this out haha, i have a 93 in my math class, the final exam is 30% of our grade whats the lowest grade that i can get on the final and still pass with an A (90)?

OpenStudy (anonymous):

93% before the exam?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

final exam is next week =(

OpenStudy (anonymous):

\[93\times\frac{70}{100}+x\times\frac{30}{100}=90\]

OpenStudy (anonymous):

93 * 0.7 = 65.1 x * 0.3 = 0.3x 0.3x + 65.1 = 90 0.3x = 24.9 x = 83 You need at least 83 percent